Monochromatic light is incident on two slits separated by 0,03 m. The distance between antinodes is 0,000128 m. If the screen is 1,6 m from the slits, what is the wavelength of light? Write your answer in scientific notation or in decimal form


1
Expert's answer
2021-06-22T11:31:07-0400

Given:

"d=0.03\:\rm m"

"\Delta y=0.000128\:\rm m"

"D=1.6\:\rm m"

The distance between antinodes

"\Delta y=\frac{D\lambda }{d}"

Hence, the wavelength of light

"\lambda=\frac{d\Delta y}{D}=\frac{(0.03\:\rm m)\times (0.000128\:\rm m)}{1.6\:\rm m}=2.4\times 10^{-6}\:\rm m"